हिंदी

In a box of floppy discs, it is known that 95% will work. A sample of three of the discs is selected at random. Find the probability that exactly one of the floppy disc work.

Advertisements
Advertisements

प्रश्न

In a box of floppy discs, it is known that 95% will work. A sample of three of the discs is selected at random. Find the probability that exactly one floppy disc work.

योग
Advertisements

उत्तर

Let X = number of working discs.

p = probability that a floppy disc works

∴ p = 95% = `95/100 = 19/20`

and q = 1 - p = `1 - 19/20 = 1/20`

Given: n = 3

∴ X ~ B`(3, 19/20)`

The p.m.f. of X is given by

P(X = x) = `"^nC_x  p^x q^(n - x)`

i.e. p(x) = `"^3C_x (19/20)^x (1/20)^(3-x)`, x = 0, 1, 2, 3

P(exactly one floppy discs work) = P(X = 1)

= p(1) = `"^3C_1 (19/20)^1 (1/20)^(3 - 1)`

`= 3 xx 19/20 xx (1/20)^2`

`= 3(19/20^3)`

Hence, the probability that none of the floppy disc will work = 3`(19/20^3)`

shaalaa.com
  क्या इस प्रश्न या उत्तर में कोई त्रुटि है?
अध्याय 8: Binomial Distribution - Exercise 8.1 [पृष्ठ २५२]

APPEARS IN

बालभारती Mathematics and Statistics 2 (Arts and Science) [English] Standard 12 Maharashtra State Board
अध्याय 8 Binomial Distribution
Exercise 8.1 | Q 9.2 | पृष्ठ २५२

वीडियो ट्यूटोरियलVIEW ALL [1]

संबंधित प्रश्न

The probability that a certain kind of component will survive a check test is 0.6. Find the probability that exactly two of the next four components tested will survive.


A die is thrown 6 times. If ‘getting an odd number’ is a success, find the probability of 5 successes. 


A die is thrown 6 times. If ‘getting an odd number’ is a success, find the probability of at least 5 successes.


A die is thrown 6 times. If ‘getting an odd number’ is a success, find the probability of at most 5 successes.


A pair of dice is thrown 4 times. If getting a doublet is considered a success, find the probability of two successes.


Five cards are drawn successively with replacement from a well-shuffled deck of 52 cards; find the probability that all the five cards are spades.


Five cards are drawn successively with replacement from a well-shuffled deck of 52 cards; find the probability that none is a spade.


In a box of floppy discs, it is known that 95% will work. A sample of three of the discs is selected at random. Find the probability that exactly two floppy disc work.


In a box of floppy discs, it is known that 95% will work. A sample of three of the discs is selected at random. Find the probability that all 3 of the sample will work.


Choose the correct option from the given alternatives:

A die is thrown 100 times. If getting an even number is considered a success, then the standard deviation of the number of successes is ______.


Choose the correct option from the given alternatives:

The mean and the variance of a binomial distribution are 4 and 2 respectively. Then the probability of 2 successes is


Choose the correct option from the given alternatives:

For a binomial distribution, n = 5. If P(X = 4) = P(X = 3), then p = ______


For a binomial distribution, n = 4. If 2P(X = 3) = 3P(X = 2), then p = ______.


If X ~ B(4, p) and P(X = 0) = `16/81`, then P(X = 4) = ______.


Choose the correct option from the given alternatives:

The probability of a shooter hitting a target is `3/4` How many minimum numbers of times must he fire so that the probability of hitting the target at least once is more than 0·99?


Let X ~ B(10, 0.2). Find P(X = 1).


Let X ~ B(10, 0.2). Find P(X ≥ 1).


Let X ~ B(10, 0.2). Find P(X ≤ 8).


The probability that a bomb will hit a target is 0.8. Find the probability that out of 10 bombs dropped, exactly 2 will miss the target.


The probability that a mountain-bike travelling along a certain track will have a tyre burst is 0.05. Find the probability that among 17 riders: at most three have a burst tyre


A large chain retailer purchases a certain kind of electronic device from a manufacturer. The manufacturer indicates that the defective rate of the device is 3%. The inspector of the retailer picks 20 items from a shipment. What is the probability that the store will receive at most one defective item?


An examination consists of 10 multiple choice questions, in each of which a candidate has to deduce which one of five suggested answers is correct. A completely unprepared student guesses each answer completely randomly. What is the probability that this student gets 8 or more questions correct? Draw the appropriate morals.


The probability that a machine will produce all bolts in a production run within specification is 0.998. A sample of 8 machines is taken at random. Calculate the probability that all 8 machines.


A computer installation has 10 terminals. Independently, the probability that any one terminal will require attention during a week is 0.1. Find the probabilities that 0.


A computer installation has 10 terminals. Independently, the probability that any one terminal will require attention during a week is 0.1. Find the probabilities that 1.


A computer installation has 10 terminals. Independently, the probability that any one terminal will require attention during a week is 0.1. Find the probabilities that 3 or more, terminals will require attention during the next week.


It is observed that it rains on 12 days out of 30 days. Find the probability that it rains exactly 3 days of week.


If E(x) > Var(x) then X follows _______.


In Binomial distribution if n is very large and probability success of p is very small such that np = m (constant) then _______ distribution is applied.


In a Binomial distribution with n = 4, if 2P(X = 3) = 3P(X = 2), then value of p is ______.


If X ~ B(n, p) with n = 10, p = 0.4, then find E(X2).


Choose the correct alternative:

A sequence of dichotomous experiments is called a sequence of Bernoulli trials if it satisfies ______


If X follows a binomial distribution with parameters n = 10 and p. If 4P(X = 6) = P(X = 4), then p = ______ 


In a binomial distribution `B(n, p = 1/4)`, if the probability of at least one success is greater than or equal to `9/10`, then n is greater than ______.


In a binomial distribution `B(n, p = 1/4)`, if the probability of at least one success is greater than or equal to `9/10`, then n is greater than ______.


If X∼B (n, p) with n = 10, p = 0.4 then E(X2) = ______.


In a binomial distribution, n = 4 and 2P(X = 3) = 3P(X = 2), then q = ______.


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×